Can some solve this problem please??
A car was moving at 12.0m/s when it hit a tree according to the damage which occurred The skid marks before the crash were 10.0m long And the acceleration must have been about -7.0 m/s^2 given the nature of the marksHow fast was the car moving before the skid?
1
Expert's answer
2019-10-07T10:47:09-0400

Let the initial speed of the car be "v_0" and the speed before hitting the tree "v = 12 \frac{m}{s}". Let the distance be "L = 10 m", and the acceleration "a = -7 \frac{m}{s^2}". The equations of the retarded motion are then:

"v = v_0 + a t", "L = v_0 t + \frac{a t^2}{2}".

Substituting "t = \frac{v-v_0}{a}" from the first equation into the second, obtain:

"L = \frac{v_0(v-v_0)}{a} + \frac{a(v-v_0)^2}{2 a^2} = \frac{1}{2 a}(v^2-v_0^2)" , from where the initial speed is "v_0 = \sqrt{v^2 - 2 S a} \approx 16.85 \frac{m}{s}".